# Marine spatial planning techniques with a case study on wave-powered offshore aquaculture farms

### v2

@@ 5.1 GIS-Python interface
The handler allows the model to query individual points by latitude and longitude and passes along the value at that point for each loaded condition raster. It then flags the point if conflicts exist with any of the conflict vectors that have been loaded. Depending on the application, more or less detail about the conflicts may be desired. In this case study, for example, one flag is used to identify conflicts, another is used to indicate whether the point queried is out of the project scope, and a third is used if all raster condition data sets exist at the given point. The handler then receives the result for a given point from the cost model and stores it in a GeoDataFrame. Although individual queries are helpful for an optimization approach, the handler can also query information for many points at once to fully populate a grid of potential sites. Finally, the GeoDataFrame with the model result is exported again as a GeoJSON file that can be loaded into QGIS to create final maps of the results.
@@ 5.2 Cost model
A metric similar to levelized cost of energy is also used to calculate the cost per fish yield of a farm at a given site. This levelized cost of fish is calculated by dividing the total farm cost by the discounted fish yield over the lifetime of the farm, as shown in equation 3 where and are the same as in equation 2.
It should also be mentioned that the cost analysis in this paper is made per an assumption that the energy demand of the aquaculture farm is modeled as a constant annual value using a reference value for power per fish yield reported by Freeman et al. (2022). Hence, the energy demand is overestimated and considered constant per year, and hourly fluctuation of the power consumption of the aquaculture farm, such as an increase in energy demand while fish feeding, is not considered in this paper. Furthermore, the uncertainty of the wave power, (i.e. the hourly data of wave significant height and wave energy period) is not considered in this paper, and all the results are presented with the annual average wave data. Therefore, the cost of the WPAF, and specifically the cost of WEC, could change in future studies using high-resolution hourly data for aquaculture farm energy demand and wave power.
@@ 7 Discussion
Overall trends include a large amount of variation between the costs of each site despite relative proximity, driven in large part by access to wave resources. The decrease in cost further offshore demonstrates that location-dependent WPAF costs are dominated by the WEC cost, as opposed to the vessel travel cost which increases with distance from shore. The optimal site at 43.7°N 68.9°W, with location-dependent costs of $30.0M is one of the furthest feasible sites from shore and has location-dependent costs several times cheaper than sites closer to land. Closer to shore, the cost increases significantly for sites in sheltered bays and other areas that would require many more WECs to power even a modest farm. This trend by far outweighs factors like the increased travel costs for deployment locations further from shore, although there are certain limits due to increasing depth and other conflicts. The main takeaway from this cost analysis is that it is worthwhile to consider deployment locations farther offshore to decrease the WPAF cost by decreasing the number of required WECs.
@@ 8 Conclusion
Marine spatial planning is a topic of critical importance given the continued development of various offshore industries and our increasing reliance on aquatic resources for food and energy. Careful planning of these resources is necessary to ensure their sustainable use and avoid conflicts with other existing and future industries that operate along our coastlines. This paper provided an overview of several MSP techniques using open source software including QGIS and Python, alongside the application of those techniques to a case study in siting a wave-powered aquaculture farm.
@@
Future work on the case study of a wave-powered aquaculture farm includes more in-depth analysis given time series data of conditions and consideration of energy storage within the WPAF to ensure a reliable power supply. Different layouts, directions, near and far-field effects, and park effects could also be integrated to provide more detailed modeling of WEC performance. The cost model could also be improved by adding terms related to the energy transfer between the WEC and aquaculture farm. An integrated multi-trophic aquaculture setup will also be considered which allows other species such as kelp and mollusks to grow alongside fish with the goal of creating mutualistic relationships between the different species. This could be coupled with an expansion of the carrying capacity model to include aspects like nutrient use and production in addition to oxygen depletion. The MSP techniques and generated map layers from this project will also be integrated into a more general web-based mapping tool to make the work more accessible to the many stakeholders who could benefit from this kind of analysis. Finally, more outreach with stakeholders will be conducted to receive feedback on the usefulness of the existing analysis and more information on the issues they face in practice.
@@ Acknowledgments
This work was supported in part by Sea Grant Regional Research Project No. R/ATD-18-NESG and the Engineering Learning Initiatives program at Cornell University. We would also like to thank Rebecca McCabe and Matthew Haefner; members of the Symbiotic Engineering and Analysis (SEA) Lab; and our industry collaborators CalWave Power Technologies, Innovasea, and Manna Fish Farms, for providing feedback and support throughout this project.
@@
Code required to reproduce the findings in this paper is available in the SEA Lab GitHub at https://github.com/symbiotic-engineering/aquaculture, and the relevant preprocessed data files are hosted on Zenodo at https://zenodo.org/records/10140826. Raw data files are publicly available from the sources cited in tables 1 and 2.
@@ Appendix A QGIS Instructions
A.1 Vector interpolation
From the processing toolbox (Processing > Toolbox), the TIN Interpolation tool can be used to convert interpolate a vector data set of point measurements into a continuous raster data set. TIN Interpolation allows the user to select the desired layer, the attribute that contains the desired measurement, and a reasonable resolution (as discussed in section 3.1) for the given project. The resulting raster is limited to the area bounded by the original points, so r.fill.stats in the processing toolbox can be used to extend the range of the raster and smooth the results of the interpolation. The parameters of this operation will vary by the project scope and individual data set. For this project, each raster is extended past the shoreline to ensure all coastal waters could be analyzed, then clipped back to include only the project scope.
A.2 Clipping
A layer used as a mask is necessary to clip layers to the desired scope. In this case, a file with state and federal waters from NOAA modified to only include the waters in the desired scope is suitable. To clip raster layers, click Raster > Extraction > Clip Raster by Mask Layer... and select the correct vector layer as a mask. For vector layers, click Vector > Geoprocessing Tools > Clip... and select the same mask as the overlay layer. Both operations will create a temporary scratch layer of the results, which can be exported to a permanent file.
A.3 Re-projection
Layers can be re-projected during the export process. Simply right-click and select Export > Save Feature As..., then select the desired projection (e.g. WGS 84) as the CRS. The CRS for final maps (e.g. WGS 84 / UTM Zone 19) is set later in the item properties of a map element in a layout, and does not need to match the CRS of individual layers.
A.4 Contour generation

### v3

@@ 5.1 GIS-Python interface
The handler allows the model to query individual points by latitude and longitude and passes along the value at that point for each loaded condition raster. It then flags the point if overlaps with any of the vectors that have been loaded. One flag is used to identify conflicts, another is used to indicate whether the point queried is out of the study scope, and a third is used if all raster condition data sets exist at the given point. The handler then receives the result for a given point from the cost model and stores it in a GeoDataFrame. Although individual queries are helpful for an optimization approach, the handler can also query information for many points at once to fully populate a grid of potential sites. Finally, the GeoDataFrame with the model result is exported again as a GeoJSON file that can be loaded into QGIS to create final maps of the results.
@@ 5.2 Cost model
A metric like levelized cost of energy is also used to calculate the cost per fish yield of a farm at a given site. This levelized cost of fish is calculated by dividing the total farm cost by the discounted fish yield over the lifetime of the farm, as shown in equation 3 where and are the same as in equation 2.
It should also be mentioned that the cost analysis in this study is made per an assumption that the energy demand of the aquaculture farm is modeled as a constant annual value using a reference value for power per fish yield reported by Freeman et al. (2022). Hence, the energy demand is overestimated and considered constant per year, and hourly fluctuation of the power consumption of the aquaculture farm, such as an increase in energy demand while fish feeding, is not considered in this study. Furthermore, the uncertainty of the wave power, (i.e. the hourly data of wave significant height and wave energy period) is not considered in this study, and all the results are presented with the annual average wave data. Therefore, the cost of the WPAF, and specifically the cost of WEC, could change in future studies using high-resolution hourly data for aquaculture farm energy demand and wave power.
@@ 7 Discussion
Overall trends include a large amount of variation between the costs of each site despite relative proximity, driven in large part by access to wave resources. The decrease in cost further offshore demonstrates that location-dependent WPAF costs are dominated by the WEC cost, as opposed to the vessel travel cost which increases with distance from shore. The optimal site at 43.7°N 68.9°W, with location-dependent costs of $30.0M is one of the furthest feasible sites from shore and has location-dependent costs several times cheaper than sites closer to land. Closer to shore, the cost increases significantly for sites in sheltered bays and other areas that would require many more WECs to power even a modest farm. This trend by far outweighs factors like the increased travel costs for deployment locations further from shore, although there are certain limits due to increasing depth and other constraints. The main takeaway from this cost analysis is that it is worthwhile to consider deployment locations farther offshore to decrease the WPAF cost by decreasing the number of required WECs. At these locations, the model estimates a levelized cost of fish near current market rates. Although this may indicate that the model is within the ballpark of actual costs, the cost model currently lacks enough detail to draw further conclusions or to make comparisons with market rates. The cost model used in this study does not include all cost components and instead focuses on those that change with location, such as distance to shore. The result is a relative, rather than absolute, metric intended to enable a comparison of different sites within the same study.
@@
Existing WPAF studies demonstrate similar methodologies and trends, although it is difficult to compare results directly given that WPAFs remain a relatively new concept. A study by Clemente et al. (2023) investigates co-location potential at two sites along the Portuguese coastline using various farm designs. Although Clemente et al. (2023) do consider some similar factors including annual wave energy availability, a point absorber WEC as an energy source, and the feasibility of Atlantic salmon, their analysis is more focused on changing the design of the farm at a limited number of sites rather than analyzing many different sites with a single farm design. A study by Garavelli et al. (2022) is more similar to this study in that it focuses on siting and considers similar geographic factors including bathymetry, wave height, and marine protected areas. Their analysis yields similar results to this study as they identify wave power density, bathymetry, and conflicts (in their case navigation routes and managed areas) to be key factors in determining favorable locations. However, they use a binary suitability score rather than a cost model making it difficult to directly compare the impact of these criteria on site selection.
@@ 8 Conclusion
Marine spatial planning is a topic of critical importance given the continued development of various offshore industries and our increasing reliance on aquatic resources for food and energy. Careful planning of these resources is necessary to ensure their sustainable use and avoid conflicts with other existing and future industries that operate along our coastlines. This paper provided an overview of several MSP techniques using open-source software including QGIS and Python, alongside the application of those techniques to a case study in siting a wave-powered aquaculture farm.
@@
Future work on the case study of a wave-powered aquaculture farm includes more in-depth analysis given time series data of conditions and consideration of energy storage within the WPAF to ensure a reliable power supply. Different layouts, directions, near and far-field effects, and park effects could also be integrated to provide more detailed modeling of WEC performance. The cost model could also be improved by adding terms related to the energy transfer between the WEC and aquaculture farm. An integrated multi-trophic aquaculture setup will also be considered which allows other species such as kelp and mollusks to grow alongside fish with the goal of creating mutualistic relationships between the different species. This could be coupled with an expansion of the carrying capacity model to include aspects like nutrient use and production in addition to oxygen depletion. Additional environmental metrics would be useful to integrate as well, including constraints on benthic habitat and areas with endangered species, and emissions savings over a diesel-powered farm. The MSP techniques and generated map layers from this study will also be integrated into a more general web-based mapping tool to make the work more accessible to the many stakeholders who could benefit from this kind of analysis. Finally, more outreach with stakeholders will be conducted to receive feedback on the usefulness of the existing analysis and more information on the issues they face in practice.
@@ Acknowledgments
This work was supported in part by Sea Grant Regional Research Project No. R/ATD-18-NESG and the Engineering Learning Initiatives program at Cornell University. We would also like to thank Rebecca McCabe and Matthew Haefner; members of the Symbiotic Engineering and Analysis (SEA) Lab; and our industry collaborators CalWave Power Technologies, Innovasea, and Manna Fish Farms, for providing feedback and support throughout this study.
@@
Code required to reproduce the findings in this study is available in the SEA Lab GitHub at https://github.com/symbiotic-engineering/aquaculture, and the relevant pre-processed data files are hosted on Zenodo at https://zenodo.org/records/10140826. Raw data files are publicly available from the sources cited in tables 1 and 2.
@@ Appendix A QGIS Instructions
A.1 File Conversion
Conversion from an ArcGIS format, is only necessary for certain raster files distributed in the proprietary GeoDatabase (.gdb) format. Although QGIS can currently access vector data in these files, rasters must be loaded into ArcGIS and exported as GeoTIFF (.tiff) files to be readable by QGIS and other applications. This process is straightforward but unfortunately requires users interested in using data from organizations that publish data in GeoDatabases to have access to ArcGIS.
A.2 Vector interpolation
Certain data sets for environmental conditions are distributed in a vector format with many points containing measurements instead of as a continuous raster. For many analyses, having a continuous raster of conditions is beneficial, and interpolation can be used to approximate this from discrete measurements. In this study, the data used for dissolved oxygen and salinity are distributed as vector files, but are converted to rasters. From the processing toolbox (Processing > Toolbox), the TIN Interpolation tool can be used to convert interpolate a vector data set of point measurements into a continuous raster data set. TIN Interpolation allows the user to select the desired layer, the attribute that contains the desired measurement, and a reasonable resolution (as discussed in section 3.1) for the given study. The resulting raster is limited to the area bounded by the original points, so r.fill.stats in the processing toolbox can be used to extend the range of the raster and smooth the results of the interpolation. The parameters of this operation will vary by the study scope and individual data set. For this study, each raster is extended past the shoreline to ensure all coastal waters could be analyzed, then clipped back to include only the study scope.
A.3 Clipping
Data is clipped to only include the scope of designated state and federal waters. This relies on a vector layer describing the study scope which, in this study, is the file with state and federal waters from NOAA in table 2. To clip raster layers, click Raster > Extraction > Clip Raster by Mask Layer... and select the correct vector layer as a mask. For vector layers, click Vector > Geoprocessing Tools > Clip... and select the same mask as the overlay layer. Both operations will create a temporary scratch layer of the results, which can be exported to a permanent file.
A.4 Re-projection
GIS data and maps use different projections to translate the three-dimensional globe into a two-dimensional image. This study uses WGS 84 (EPSG:4326), a very common CRS, to standardize all downloaded data because of its simplicity, and so that each data set uses units of degrees in latitude and longitude. WGS 84 / UTM Zone 19 is used for final maps exported from QGIS, which is tailored for the specific region of the earth included in the scope and minimizes distortion in that area. Layers can be re-projected during the export process. Simply right-click and select Export > Save Feature As..., then select the desired projection (e.g. WGS 84) as the CRS. The CRS for final maps (e.g. WGS 84 / UTM Zone 19) is set later in the item properties of a map element in a layout, and does not need to match the CRS of individual layers.
A.5 Contour generation